Given two events A and B such that P(A) = 0.67, P(B)= 0.23 and P (A /intersection/ and B) = 0.12, find each probability.
A. P(A')
B. P(B')
C. P(A U B)
D. P(A U B)'

Answer by ikleyn(52878)

.

Given two events A and B such that P(A) = 0.67, P(B)= 0.23 and P (A /intersection/ and B) = 0.12, find each probability.

A.  P(A')     = 1 - P(A) = 1 - 0.67 = 0.33.


B.  P(B')     = 1 - P(B) = 1 - 0.23 = 0.77.


C.  P(A U B)  = P(A) + P(B) - P(A intersection B) = 0.67 + 0.23 - 0.12 = 0.78.   <<<---=== it is VERY important formula,

                                               and those who know it, have a huge advantage comparing with those who do not know (!)


D.  P(A U B)' = 1 - P(A U B) = 1 - 0.78 = 0.22.

Solved.   //   All questions are answered.